Balls of the same size and weight are placed in a container. There are 8 different colours and 90 balls in each colour.
What is the minimum number of balls that much be drawn from the container in order to get balls of 4 different colours with at least 9 balls for each colour.

Answer by greenestamps(13200)



The worst case scenario is that you get all 90 balls of 3 of the colors and 8 balls each of the other 5 colors. After that, the next ball you draw has to give you at least 9 balls of 4 different colors.

ANSWER: 3(90)+5(8)+1 = 270+40+1 = 311